OpenStudy (anonymous):

change the equation into slope-intercept form. And Identify the slope and y-intercept of the equation. –3x + 2y = 8

OpenStudy (anonymous):

Find x and y intercept

OpenStudy (angela210793):

solve for y y=mx+b where m will be the slope and b will be y-intercept

OpenStudy (anonymous):

\[\frac{ x }{ a}+\frac{ y }{ b}=1\]

OpenStudy (anonymous):

for y intercept put x=0 in the equation and vice versa